OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (anonymous):

13.

OpenStudy (asnaseer):

what you need to here is express the problem using algebra. so we have: 1. Coach Johnson is 10 years older than Tim. Lets call Coach Johnsons age "J" and Tims age "T". So J=10+T 2. When Tim started football 3 years ago, Coach Johnson was twice his age. So 5 years ago, Coach Johnson was (J-5) years old and Tim was (T-5) years old. We therefore get (J-5)=2(T-5) so you have these equations: J = 10 + T J-5 = 2(T-5) 2 equations and 2 unknowns which you should be able to solve.

OpenStudy (asnaseer):

sorry I mis-read your question, it should be 3 years ago that Tim started football NOT 5 years ago, so the equations are: J = 10 + T J-3 = 2(T-3)

OpenStudy (asnaseer):

if you then place the value of J from the 1st equation into the 2nd you will get: (10+T)-3 = 2(T-3) 10+T-3=2T-6 7+T=2T-6 7+6=2T-T 13=T

OpenStudy (asnaseer):

I hope you understand how this was achieved?